An airport terminal. The upper level is 4.5 m above the lower level, and the length of the escalator is 7.6 m . Assume that your mass is 66 kg.

How much work does the up escalator do on you when you ride it from the lower level to the upper level?

How much work does the down escalator do on you when you ride it from the upper level to the lower level?
Answer #1

by the energy coservation law work done by escalator on you = mgh

where m is mass

g is garvity

h is height from lower level to upper level

so in first part

in work done the up escalator do on you = +mgh = 66*9.8*4.5= 2910.6J

in second part

escalator going down mean work don on escaltor so work done is negative = -mgh=-66*9.8*4.5= -2910.6J